Estimate the values of c that satisfy the conclusion of the Mean Value Theorem on the interval (0,8). Enter your answer as a comma-separated list. Round to one decimal.


1
Expert's answer
2022-03-15T15:13:24-0400

Solution


According to the mean value theorem,


If  "f(x)" is continuous on interval "[a, b]" , and differentiable on "(a, b)" .


Then there exists "c" such that "a < c <b" and 


"f'\left( c \right) = \frac{{f\left( b \right) - f\left( a \right)}}{{b - a}}"


In the question the function is missing, so let us choose, the function as


"f\left( x \right) = {x^3} + 2x - 3" which is continuous with in the interval "[0,8]" and differentiable within "(0, 8)" ,

then for "0 < c <8", we have


"f'\left( c \right) = \frac{{f\left( 8 \right) - f\left( 0 \right)}}{{8 - 0}}"


"f'\left( c \right) = \frac{{\left[ {{{\left( 0 \right)}^3} - 2\left( 0 \right) - 3} \right] - \left[ {{{\left( 8 \right)}^3} - 2\left( 8 \right) - 3} \right]}}{{8 - 0}}\\"


"f'\left( c \right) = \frac{{ - 3 - 493}}{8}\\"


"f'\left( c \right) = 62"


Now


"f\left( x \right) = {x^3} + 2x - 3\\"


"f'\left( x \right) = 3{x^2} + 2\\"


"f'\left( c \right) = 3{c^2} + 2"


Therefore,


"3{c^2} + 2=62"


"3{c^2} =62-2"


"c^2=\frac{60}{3}=20"


"c=2\sqrt 5 \" (ignoring negative value since we need to focus within the interval "[0, 8]" )


since we can see that


"0<2\sqrt 5 < 8"


Hence required value of "c" is "2\sqrt 5"